Grade 11Physical Chemistry

6.3g of oxalic acid dihydrate have been dissolved in water to obtain 250ml solution. How much volume of 0.1N NaOH would be required to neutralize 10ml of this solution?

Soumya Ranjan Swain
9 Years agoGrade 11

1 Answer

Khushi goyal
9 Years ago
Gm eqv of H2C2O4.2H2O IN 250 mL = 6.3/126×2 =0.1 In 10 ml = 0.1×10 /250 =0.004Using euivalence concept gm Equivalence of H2C204.2H20= gm eqv. Of NaoH0.004=0.1×V (in L)V=0.04 L=40 mL Hope it helps :) If then Click approved.